OpenStudy (nautika27):

The perimeter of a rectangle is 38 inches. If the length of the rectangle is 14 inches, which equation could be used to find the width, x? A. x + 2(14) = 38 B. 2(x + 14) = 38 C. 14(x + 2) = 38 D. x + 14 = 38

OpenStudy (nautika27):

@Kanard

jhonyy9 (jhonyy9):

do you know that the perimeter of a rectangle is 2w+2l yes ?

jhonyy9 (jhonyy9):

than is it given that P=38 and l=14 so how you can rewrite it ?

OpenStudy (nautika27):

c

jhonyy9 (jhonyy9):

so used the width sign x

OpenStudy (nautika27):

huh ?

OpenStudy (nautika27):

was i wrong ?

jhonyy9 (jhonyy9):

38=2*14 +2x 38=2(14+x) so than what is right ?

OpenStudy (nautika27):

b

jhonyy9 (jhonyy9):

yes b. is right sure hope you understand it now yes ?

OpenStudy (nautika27):

yes thank you do you think you could help me with one more just one

jhonyy9 (jhonyy9):

yes sure but please write it on the open question column

jhonyy9 (jhonyy9):

and tag me there